# Semiparametric Preference Optimization: Your Language Model is Secretly a Single-Index Model

### v1

@@ Abstract
Aligning large language models to preference data is commonly implemented by assuming a known link function between the distribution of observed preferences and the unobserved rewards (e.g., a logistic link as in Bradley-Terry). If the link is wrong, however, inferred rewards can be biased and policies be misaligned. We study policy alignment to preferences under an unknown and unrestricted link. We consider an $f$-divergence-constrained reward maximization problem and show that realizability of the solution in a policy class implies a semiparametric single-index binary choice model, where a scalar-valued index determined by a policy captures the dependence on demonstrations and the rest of the preference distribution is an unrestricted function thereof. Rather than focus on estimation of identifiable finite-dimensional structural parameters in the index as in econometrics, we focus on policy learning, focusing on error to the optimal policy and allowing unidentifiable and nonparametric indices. We develop a variety of policy learners based on profiling the link function, orthogonalizing the link function, and using link-agnostic bipartite ranking objectives. We analyze these and provide finite-sample policy error bounds that depend on generic functional complexity measures of the index class. We further consider practical implementations using first-order optimization suited to neural networks and batched data. The resulting methods are robust to unknown preference noise distribution and scale, while preserving the direct optimization of policies without explicitly fitting rewards.
@@ 1 Introduction
In this paper, we study instead LLM alignment from preference data generated by a completely unknown link function and unknown reward function, where the target is the optimal reward-maximizing divergence-constrained policy. A key insight is that mere realizability of this optimal policy in our LLM policy class implies a so-called semiparametric single-index model for preferences, which has an arbitrary unknown link function. The policy class induces a class of indices (corresponding to reward or utility functions), which can be an arbitrary nonparametric class, and moreover the true index (corresponding to the true reward function) is generally not identifiable. We study policy learning methods and analyze bounds on the policy difference to the optimal policy (rather than convergence of structural estimates). Finally, we give practical gradient-based optimization algorithms that can be used to align neural-net-based models such as transformer LLMs.
We call the resulting general approach Semiparametric Preference Optimization (SPO). The primary algorithm we study is Orthogonalized SPO (OSPO), which performs empirical risk minimization with a loss function that depends on a family of univariate regression functions related to the unknown link. The analysis shows we can simply plug in estimates for these regression (for example, kernel regression estimates, since the regression is just univariate) and the method is locally invariant to the estimation: the error bounds scale quadratically in the error of the plug-in. This shows how we can perform almost as well as if we had known the link exactly (as, for example, DPO assumes we do) while being robust to not knowing the link at all (e.g., even when it is not logistic). In summary, SPO keeps the divergence-constrained policy search without an explicit reward model of DPO, while removing unverifiable assumptions on the reward-to-preference link and still provably optimally aligning policies to preferences.
@@
  • From RLHF to single-index model. We formulate realizability of the $f$-divergence-constrained optimal policy and preferences observed with unknown link function as a single-index model of binary choice, with an index given by our policy class.
  • Semiparametric policy learners for RLHF. We develop the policy learners Profiled SPO (PSPO), Orthogonalized SPO (OSPO), Ranking SPO (RSPO) as three ways to search over policies that would maximize the likelihood or rank of their implied index in this model, without explicitly learning a reward model.
  • Policy-learning theory. We formalize an affine-invariant metric on indices and prove it controls policy difference to the optimal policy. We analyze guarantees of our policy learners in terms of this metric, showing consistency of PSPO (without a rate) and establishing rates for OSPO and RSPO that naturally depend on the functional complexity of the given classes.
  • Practical algorithms. We provide practical first-order algorithms suited to neural networks and batched data and study a simple empirical recalibration step at the end that enforces the divergence budget.
@@
Section 2 derives the single-index formulation and the implications for policy learning. Section 3 presents and analyzes PSPO. Section 4 presents and analyzes OSPO. Section 5 presents and analyzes RSPO. Section 6 analyzes empirical calibration to the divergence budget. Section 7 discusses possible extensions. Section 8 surveys related work. Full proofs appear in the appendices.
@@ 2.3 Reformulation as a single-index model
$h_{\theta}(x,y)$ $=f^{\prime}\left(\frac{\pi_{\theta}(y\mid x)}{\pi_{\rm ref}(y\mid x)}\right),\qquad t_{\theta}(x,y_{0},y_{1})=h_{\theta}(x,y_{1})-h_{\theta}(x,y_{0}).$ (5)
@@
may be parametric or nonparametric and is in general as rich as the policy class $\{\pi_{\theta}:\theta\in\Theta\}$. As such, we will characterize $\mathcal{T}$ in terms of its own functional complexity, permitting rich and growing function classes.
@@ 3.3 First-order optimization
To optimize Eq. 6 in practice, especially when $\pi_{\theta}$ is a neural network, we can use a first-order (gradient-based) method. A simple approach would be to alternate taking a gradient step in $\theta$, taking $\Psi$ as fixed, with optimizing $\Psi$ for the current $\theta$. The former can be combined with any momentum and/or adaptive learning rate procedure. The latter is essentially a univariate monotone logistic regression and can be done using the Pool Adjacent Violators Algorithm (PAVA). This is akin to the Isotron algorithm (Kalai and Sastry, 2009; Kakade et al., 2011). Alternatively, differentiation through the isotonic regression can be done by implicit differentiation of PAVA KKT conditions or by training a separate network for $\Psi$ with monotonic constraints.
@@ 5.3 First-order optimization
A more appealing alternative is to use a surrogate loss function for bipartite ranking (Agarwal, 2014). In particular, we can obtain one from any surrogate loss for classification (Bartlett et al., 2006) by transforming the data into classification data. Let us focus on transforming Eq. 13, as that is the variant we recommend in practice; the transformation is analogous for Eq. 12. Create the $2n^{2}$ data points, where for each $i,j$, we label the features $(x_{j},y_{j,1-z_{j}},y_{j,z_{j}},x_{i},y_{i,z_{i}},y_{i,1-z_{i}})$ with $1$ and the features $(x_{j},y_{j,z_{j}},y_{j,1-z_{j}},x_{i},y_{i,1-z_{i}},y_{i,z_{i}})$ with 0. Then Eq. 13 is given by $2$ times the 0-1 loss of this classification problem over class scores $(w,w^{\prime})\mapsto t(w)-t(w^{\prime})$. Now use any convex classification surrogate loss, such as hinge or logistic. Now, rather than taking batches over the $n^{2}$ augmented data points, we can take batches of size $B$ from the original $n$ data points and simply consider the gradient for the corresponding $B^{2}$ terms.
@@ 9 Conclusion
Divergence-constrained alignment of LLMs naturally induces a semiparametric single-index model with the LLM policy class parametrizing the index. This requires only the probability one prefers one demonstration over another be some unknown increasing function of the difference in rewards. This significantly relaxes assuming the link function is known exactly, as in DPO. However, it means we no longer have a clean likelihood in the policy parameter alone. Inspired by Cosslett (1983), we considered an orthogonalized quasi-likelihood with a plug-in for a high-dimensional nuisance function characterizing the distribution of preference given any candidate index. Our analysis gave finite-sample bounds and admitted nonparametric function classes with indices identifiable only up to scale. The dependence on plug-in error was squared, meaning it generally does not affect the learning rate, and the learner behaves as though the link function were known. The semiparametric preference optimization framework developed thus allows to align LLMs to human and AI feedback in a way robust to how these are linked to rewards.
@@ D.4.2 Proof of the theorem
Let $\delta\in(0,0.5)$. Intersect the events of Lemma 16 (at confidence $\delta/3$) and Lemma 18 (at confidence $\delta/3$). On this intersection, Eq. 14 holds and the bounds Eqs. 15 and 16 hold.
@@
Apply Lemma 21, then substitute Eqs. 15 and 16 and $b_{+}\asymp h$, $b_{-}\asymp h$ from Eq. 14. This yields
@@ E.2 Proof of Theorem 7
Furthermore, for any measurable $g:\mathbb{R}\to\mathbb{R}$, the score $g(t)$ is measurable w.r.t. $\sigma(t)$, and by Eq. 20, the conditional expectation given $(t_{1},t_{0})$ is maximized by choosing the sign $\mathrm{sgn}_{0}(g(t_{1})-g(t_{0}))$ equal to $\mathrm{sgn}_{0}(\bar{\eta}_{1}-\bar{\eta}_{0})$. Hence $\mathrm{AUC}(g(t))\leq\mathrm{AUC}(\bar{\eta})$ for all $g$, in particular $\mathrm{AUC}(t)\leq\mathrm{AUC}(\bar{\eta})$. Combining Eqs. 21 and 22 and $\mathrm{AUC}(t)\leq\mathrm{AUC}(\bar{\eta})$ yields
@@
Under (H2)–(H5), $\eta(w)=\Psi^{\star}(t^{\star}(w))$ has support an interval $J=[\Psi^{\star}(\underline{t^{\star}}),\Psi^{\star}(\overline{t^{\star}})]$ and admits a density $q$ given by the change-of-variables formula $q(s)=p(r)/(\Psi^{\star})^{\prime}(r)$ where $s=\Psi^{\star}(r)$. Thus $q(s)\geq\underline{p}/\overline{\psi}=\underline{q}$ on $J$. Applying Lemma 24 with $U=\eta(w)$ and $\mathcal{F}=\sigma(t(w))$ gives
@@
Plugging into Eq. 27 and using (H6) yields the global inequality
@@
We need a continuity bound comparing $\mathrm{AUC}(s)$ and $\mathrm{AUC}(-t^{\star})$. First, under (H2)–(H4), the conditional density of $r=t^{\star}(w)$ given $z=1$ is $p_{1}(r)=\eta(r)p(r)/\pi_{1}$, hence

### v2

@@ Abstract
Aligning large language models (LLMs) to preference data typically assumes a known link function between observed preferences and latent rewards (e.g., a logistic Bradley-Terry link). Misspecification of this link can bias inferred rewards and misalign learned policies. We study preference alignment under an unknown and unrestricted link function. We show that realizability of $f$-divergence-constrained reward maximization in a policy class induces a semiparametric single-index binary choice model, where a scalar policy-dependent index captures all dependence on demonstrations and the remaining preference distribution is unrestricted. Rather than assuming this model has identifiable finite-dimensional structural parameters and estimating them, as in econometrics, we focus on policy learning with the reward function implicit, analyzing error to the optimal policy and allowing for unidentifiable nonparametric indices. We develop preference optimization algorithms robust to the unknown link and prove convergence guarantees in terms of generic function complexity measures. We demonstrate this empirically on LLM alignment. Code is available at https://github.com/causalml/spo/.
@@ 1 Introduction
In this paper, we study instead LLM alignment from preference data generated by a completely unknown link function and unknown reward function, where the target is the optimal reward-maximizing divergence-constrained policy. A key insight is that mere realizability of this optimal policy in our LLM policy class implies a so-called semiparametric single-index model for preferences, which has an arbitrary unknown link function. The policy class induces a class of indices (corresponding to reward or utility functions), which can be an arbitrary nonparametric class, and moreover the true index (corresponding to the true reward function) is generally not identifiable. We study policy learning methods and analyze bounds on the policy difference to the optimal policy. We call the resulting approach Semiparametric Preference Optimization (SPO).
@@
  • From RLHF to a single-index model. We show that realizability of the $f$-divergence-constrained optimal policy with unknown link induces a semiparametric single-index choice model, with an index determined by the policy class. We show the optimal policy is generally learnable even thought the index is not.
  • Semiparametric policy learners. We develop the policy learners Profiled SPO (PSPO), Orthogonalized SPO (OSPO), and Ranking SPO (RSPO) as three ways to optimize over policies without specifying the link or explicitly fitting a reward model. PSPO maximizes the likelihood also over the unknown link and provides a proof of concept that specifying the link can be avoided for policy learning. OSPO leverages a plug-in estimate for a nuisance regression related to the link function and to first order behaves as though the link function were known. RSPO avoids nuisance estimation and, while it can be less statistically efficient, it can be instantiated as a simple and practical “pairs-of-pairs” variant of DPO (PoP-DPO) that can provide additional robustness to the link function.
  • Policy-learning theory. We establish policy convergence guarantees for our methods under functional complexity conditions that admit generic policy classes (and implied reward index classes). These leverage analyses of the statistics of each empirical optimization problem, how near optimality translates to index error up to its non-identifiable aspects, how that in turn translates to policy error, and how plug-in nuisance error can have vanishing impact on performance.
  • Algorithms and experiments. We provide practical first-order algorithms suited to neural networks and batched data and an empirical recalibration step to enforce the divergence budget. We demonstrate these methods numerically first in an illustrative synthetic example and then in LLM alignment using the Qwen3 model and UltraFeedback dataset.
@@
Section 2 derives the single-index formulation and the implications for policy learning. Section 3 presents and analyzes PSPO. Section 4 presents and analyzes OSPO. Section 5 presents and analyzes RSPO. Section 6 analyzes empirical calibration to the divergence budget. Section 7 provides an empirical investigation of the algorithms, including in LLM alignment. Section 8 discusses possible extensions. Section 9 surveys related work. Full proofs appear in the appendices.
@@ 2.3 Reformulation as a single-index model
$h_{\theta}(x,y)$ $=f^{\prime}\left(\frac{\pi_{\theta}(y\mid x)}{\pi_{\rm ref}(y\mid x)}\right),$ (5)
$t_{\theta}(x,y_{0},y_{1})$ $=h_{\theta}(x,y_{1})-h_{\theta}(x,y_{0}).$ (6)
@@
may be parametric or nonparametric and is in general as rich as the policy class $\{\pi_{\theta}:\theta\in\Theta\}$. As such, we will characterize $\mathcal{T}$ in terms of its own functional complexity, permitting rich classes and treating them generically.
@@ 3.3 First-order optimization
To optimize Eq. 7 in practice, especially when $\pi_{\theta}$ is a neural network, we can use a first-order (gradient-based) method. A simple approach would be to alternate between taking a (batched stochastic) gradient step in $\theta$, taking $\Psi$ as fixed, and optimizing $\Psi$ for the current $\theta$. The former can be combined with any momentum and/or adaptive learning rate procedure. The latter is essentially a univariate monotone logistic regression and can be done using the Pool Adjacent Violators Algorithm (PAVA). This is akin to the Isotron algorithm (Kalai and Sastry, 2009; Kakade et al., 2011). Alternatively, differentiation through the isotonic regression can be done by implicit differentiation of PAVA KKT conditions or by training a separate network for $\Psi$ with monotonic constraints.
@@ 4.7 First-order optimization
Given a differentiable $\hat{g}_{\theta}$ (e.g., kernel regression), we can directly apply standard first-order optimizers to Eq. 9.
@@ 5.1 From AUC maximization to an index and a policy
Note that unlike Lemma 1, here we can restrict to nonnegative scalings. Indeed, AUC is not invariant to flipping the sign – that will yield the complement of the AUC. Our regularity conditions above ensure that the optimal AUC (attained by $t^{\star}$) is strictly above $0.5$ and that the AUC is $\frac{2}{3}$-Hölder continuous in the score function. Since AUC is invariant to positive scalings, this means that for $\iota$ small enough, the covariance of $t^{\star}$ and $t_{\theta}$ cannot be negative.
@@ 5.3 First-order optimization
A more appealing alternative is to use a surrogate loss function for bipartite ranking (Agarwal, 2014). In particular, we can obtain one from any surrogate loss for classification (Bartlett et al., 2006) by transforming the data into classification data. Let us focus on transforming Eq. 14, as that is the variant we recommend in practice; the transformation is analogous for Eq. 13. Create the $2n^{2}$ data points, where for each $i,j$, we label the features $(x_{j},y_{j,1-z_{j}},y_{j,z_{j}},x_{i},y_{i,z_{i}},y_{i,1-z_{i}})$ with 11 and the features $(x_{j},y_{j,z_{j}},y_{j,1-z_{j}},x_{i},y_{i,1-z_{i}},y_{i,z_{i}})$ with 0. Then Eq. 14 is given by $2$ times the 0-1 loss of this classification problem over class scores $(w,w^{\prime})\mapsto t(w)-t(w^{\prime})$. Now use any convex classification surrogate loss, such as hinge or logistic. To use batched first-order optimization, rather than taking batches over the $n^{2}$ augmented data points, we can take batches of size $B$ from the original $n$ data points and simply consider the gradient for the corresponding $B^{2}$ terms, as such would still form a stochastic gradient.
Remark 4 (Pairs-of-Pairs DPO).
Using the symmetric-link AUC objective in Eq. 14, replacing $\phi$ with a logistic surrogate loss ($\log(1+e^{-u})$), and instantiating with KL divergence ($f(u)=u\log u$) yields a pairs-of-pairs DPO (PoP-DPO) variant:
$\hat{\theta}_{\rm PoP-DPO}\in\operatornamewithlimits{argmin}_{\theta\in\Theta}\frac{1}{n^{2}}\sum_{i,j}\log\bigl(1+\frac{\pi_{\theta}(y_{j,1-z_{j}}\mid x_{j})}{\pi_{\rm ref}(y_{j,1-z_{j}}\mid x_{j})}\frac{\pi_{\theta}(y_{i,1-z_{i}}\mid x_{i})}{\pi_{\rm ref}(y_{i,1-z_{i}}\mid x_{i})}\frac{\pi_{\rm ref}(y_{j,z_{j}}\mid x_{j})}{\pi_{\theta}(y_{j,z_{j}}\mid x_{j})}\frac{\pi_{\rm ref}(y_{j,z_{j}}\mid x_{j})}{\pi_{\theta}(y_{j,z_{j}}\mid x_{j})}\bigr).$
This is similar to DPO’s objective, minimizing $\frac{1}{n}\sum_{i}\log\left(1+\frac{\pi_{\theta}(y_{i,1-z_{i}}\mid x_{i})}{\pi_{\rm ref}(y_{i,1-z_{i}}\mid x_{i})}\frac{\pi_{\rm ref}(y_{i,z_{i}}\mid x_{i})}{\pi_{\theta}(y_{i,z_{i}}\mid x_{i})}\right)$. Seen as a surrogate-loss classification algorithm, DPO seeks consistent ranking within each demonstration pair, while PoP-DPO compares pairs of pairs and is able to explore a greater range of contrasts rather than rely on the link to extrapolate from just within-pair comparisons, offering some additional robustness, since we are not interested in simply the classification quality but in the fidelity of the learned reward/index.
@@ 6 Empirical Recalibration to the Divergence Budget
Remark 5 (Sampling from $\pi_{\beta,\theta}(y\mid x)$ with autoregressive LLMs).
In the case of KL divergence ($f(u)=u\log u$), we have $\pi_{\beta,\theta}(y\mid x)\propto(\pi_{\rm ref}(y\mid x))^{1-1/\beta}(\pi_{\theta}(y\mid x))^{1/\beta}$. If $y=(y_{1},y_{2},\dots)$ consists of a sequence of tokens and $\pi_{\rm ref}(y\mid x)$, $\pi_{\theta}(y\mid x)$ are given in the form of token-by-token conditionals $\pi_{\rm ref}(y_{t}\mid x,y_{<t})$, $\pi_{\theta}(y_{t}\mid x,y_{<t})$, which are in turn given by a softmax distribution over some (not necessarily normalized) logits $g_{\rm ref}(y_{t}\mid x,y_{<t})$, $g_{\theta}(y_{t}\mid x,y_{<t})$, then to sample from $\pi_{\beta,\theta}(y\mid x)$ we can simply sample token-by-token from the softmax distribution over the logits $(1-1/\beta)g_{\rm ref}(y_{t}\mid x,y_{<t})+(1/\beta)g_{\theta}(y_{t}\mid x,y_{<t})$.
7 Empirical investigation
[…]
7.1 Synthetic preference optimization experiment
First, we consider an illustrative small synthetic setting. We simulate preference data with 20-dimensional Gaussian contexts $x$ and $\left\lvert\mathcal{Y}\right\rvert=10$ actions. We set the reference policy to uniform and consider a policy class $\pi_{\theta}$ given by a feedforward network with two hidden layers. We fix some $\theta^{\star}$ and set $r^{\star}(x,y)=10\log(\pi_{\theta^{\star}}(y\mid x)/\pi_{\mathrm{ref}}(y\mid x))$. Preferences are generated from an even mixture of two shifted logistic links $0.5\,\sigma(4(u-s))+0.5\,\sigma(4(u+s))$ with shift $s$ controlling misspecification ($s=0$ corresponds to logistic link). Each method learns $\pi_{\theta}$ from $n=1000$ preference pairs, and we repeat the experiment 1000 times. Additional detail is given in Section G.1.
(a) Reward at a given KL budget.
(b) Achievable reward-KL, $s=0$.
(c) Achievable reward-KL, $s=1.5$.
Figure 1: Synthetic preference optimization experiment.
Figure 1 shows the average reward (and 90% confidence intervals) at KL budget $\kappa=0.2$ as $s\in[0,1.5]$ varies and the average reward-divergence curves for $\pi_{\beta,\hat{\theta}}$ as $\beta$ varies, at $s=0$ and $s=1.5$. All methods perform well when logistic is well-specified ($s=0$). DPO performance deteriorates as $s$ grows, and so does RSPO due to its logistic surrogate loss, albeit providing some robustness at low misspecification. PSPO performs less well, consistent with its rough loss surface, but is robust to misspecified link, with performance roughly flat as $s$ varies. There is some light deterioration due to the lower SNR given the link is very flat around 0 when $s$ is large (rather than due to misspecification). OSPO consistently performs the best across all $s$, consistent with it being efficient no matter what the link is, whether logistic or not.
7.2 Aligning Qwen3 on UltraFeedback
We next consider an experiment aligning the Qwen3 model (Yang et al., 2025) using the UltraFeedback data set (Cui et al., 2024) and Skywork-Reward-V2 reward model (Liu et al., 2025). We take as our reference model Qwen3-0.6B after some light supervised fine-tuning on UltraFeedback responses to ensure format adherence. In order to be able to evaluate and compare learned policies, we take Skywork-Reward-V2-Qwen3-1.7B as the ground truth reward function. For prompts from UltraFeedback, we generate two responses from the reference policy, score them with the reward model, and take the difference, normalizing so that differences have variance 1 across the data. We generate preference observations using the same mixture of shifted logitic links as in the last section. We run each method on the preference observations alone (no reward) to train a rank-16 LoRA. We then vary $\beta$, use Remark 5 to generate responses from the aligned and scaled policy, and score using the reward model. Additional detail is given in Section G.2.
Figure 2 depicts the resulting average reward over 4 replications at $\kappa=0.5$ (with 90% confidence intervals for the difference to DPO). It overall appears to convey a similar story as the synthetic experiment, albeit with more noise, given only four replications. We again see all methods perform similarly with a well-specified logistic link, DPO deteriorating when the logistic link becomes more misspecified, and the SPO methods remaining robust to a non-logistic link. OSPO and PSPO seem to be noisier and possibly unstable, consistent with their more intricate training procedures. RSPO (as PoP-DPO) seems to strike a good balance of simplicity and robustness, offering a practical approach to still use logistic loss optimization while squeezing more comparisons from preference data rather than rely on a well-specified link to extrapolate, given we are interested in more than just classification.
Figure 2: Aligning Qwen3 on UltraFeedback.
@@ 10 Conclusion
Divergence-constrained alignment of LLMs induces a semiparametric single-index model with the policy class parametrizing the index. This significantly relaxes assuming the link function is known exactly, as in DPO. We considered policy learning by maximizing the profiled likelihood, orthogonolized quasi-likelihood, or bipartite ranking accuracy of the index implied by the policy. Our analysis gave finite-sample bounds and admitted nonparametric indices identifiable only up to scale. Empirical demonstrations showed how the SPO framework developed allows efficient alignment of LLMs to human and AI feedback in a way robust to how these are linked to rewards.
@@ D.4.2 Proof of the theorem
Let $\delta\in(0,0.5)$. Intersect the events of Lemma 16 (at confidence $\delta/3$) and Lemma 18 (at confidence $\delta/3$). On this intersection, Eq. 15 holds and the bounds Eqs. 16 and 17 hold.
@@
Apply Lemma 21, then substitute Eqs. 16 and 17 and $b_{+}\asymp h$, $b_{-}\asymp h$ from Eq. 15. This yields
@@ E.2 Proof of Theorem 7
Furthermore, for any measurable $g:\mathbb{R}\to\mathbb{R}$, the score $g(t)$ is measurable w.r.t. $\sigma(t)$, and by Eq. 21, the conditional expectation given $(t_{1},t_{0})$ is maximized by choosing the sign $\mathrm{sgn}_{0}(g(t_{1})-g(t_{0}))$ equal to $\mathrm{sgn}_{0}(\bar{\eta}_{1}-\bar{\eta}_{0})$. Hence $\mathrm{AUC}(g(t))\leq\mathrm{AUC}(\bar{\eta})$ for all $g$, in particular $\mathrm{AUC}(t)\leq\mathrm{AUC}(\bar{\eta})$. Combining Eqs. 22 and 23 and $\mathrm{AUC}(t)\leq\mathrm{AUC}(\bar{\eta})$ yields
@@
Under our assumptions, $\eta(w)=\Psi^{\star}(t^{\star}(w))$ has support an interval $J=[\Psi^{\star}(\underline{t^{\star}}),\Psi^{\star}(\overline{t^{\star}})]$ and admits a density $q$ given by the change-of-variables formula $q(s)=p(r)/(\Psi^{\star})^{\prime}(r)$ where $s=\Psi^{\star}(r)$. Thus $q(s)\geq\underline{p}/\overline{\psi}=\underline{q}$ on $J$. Applying Lemma 24 with $U=\eta(w)$ and $\mathcal{F}=\sigma(t(w))$ gives
@@
Plugging into Eq. 28 and using Assumption 5 yields the global inequality
@@
We need a continuity bound comparing $\mathrm{AUC}(s)$ and $\mathrm{AUC}(-t^{\star})$. First, the conditional density of $r=t^{\star}(w)$ given $z=1$ is $p_{1}(r)=\eta(r)p(r)/\pi_{1}$, hence